Solve by the Elimination method 2a+3b=7, -2a+b=13
Answer by checkley79(3341)

2a+3b=7
-2a+b=13 add
------------------
4b=20
b=20/4
b=5 ans.
2a+3*5=7
2a+15=7
2a=7-15
2a=-8
a=-8/2
a=-4 ans.
Proof:
-2*-4+5=13
8+5=13
13=13